Show that if the room temperature changes by a small amount from T to T + ∆T, the fundamental frequency of an organ pipe changes from v to v + ∆v, where ∆v/v = 1/2 ∆T/T.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. V base rod = ?, V base air = 340 m/s, L base r = 25 × 10^–2, d base 2 = 5 × 10^–2 metres V base r/V base a ⇒ V base r = 340 * 25 * 10^-2 * 2/5 * 10^-2 = 3400 m/s.